In your classroom, there are 8 white chairs.There are 2 times many as yellow chairs as white chairs, and the blue chairs are 3 fewer than yellow chairs.By means of a calculations determine how many chairs are in your class.
1
Expert's answer
2020-06-07T14:55:54-0400

Given that,

Number of white chairs = 8.

Number of yellow chairs = 2*8 = 16 chairs.

Number of blue chairs = Number of yellow chairs - 3 = 16 - 3 =13.

Therefore, total number of chairs in the class = 8+16+13 = 37.